OpenStudy (anonymous):

The height of a cylinder is 3 times larger than the radius. Find the volume of the cylinder as a function of the height. (V = πr^2h)

OpenStudy (johnweldon1993):

Same thing here...we have \[\large V = \pi r^2 h\] We know that the height is 3 times larger than the radius...so \[\large h = 3r\] This means that the radius is 3 times smaller than the height...so we have \[\large r = \frac{h}{3}\] Now we can replace 'r' in our first equation here...with the 'h/3' \[\large V = \pi (\frac{h}{3})^2h\] And simplify from there

OpenStudy (anonymous):

would it be V=1/3 πr^2?

OpenStudy (anonymous):

i mean to the third power

OpenStudy (johnweldon1993):

Not quite...it should be a function of 'height' right? so the only variable should be 'h' in there right?

OpenStudy (anonymous):

lol yes sorry my mistake

OpenStudy (johnweldon1993):

We have \[\large V = \pi(\frac{h}{3})^2h\] Remember that \[\large (\frac{h}{3})^2 = \frac{h^2}{3^2} = \frac{h^2}{9}\] so we have \[\large V = \pi \frac{h^2}{9}h \] \[\large V = \pi \frac{h^3}{9}\]

OpenStudy (anonymous):

V = 3πh^3 V = 9πh^3 V =1/9 πh^3 V =1/3πh^3

OpenStudy (johnweldon1993):

mmhmm same thing as \[\large V = \frac{1}{9}\pi h^3\]:)

OpenStudy (anonymous):

ohhhh okay lol thanks again ;*

OpenStudy (johnweldon1993):

Haha you're welcome again ;)